Skysong Company purchases equipment on January 1, Year 1, at a cost of $504,000. The asset is expected to have a service lifeX Your answer is incorrect. Compute the amount of depreciation for each of Years 1 through 3 using the double-declining-balanhelp!
Answer #1

Solution:

DDB rate = (1/12) * 2 = 16.67%

Depreciation Year 1 = $504000*16.67% = $84017

Depreciation for Year 2 = Book value * 16.67% = ($504000 - $84017) * 16.67% = $70011

Depreciation for Year 3 = Book value * 16.67% = ($504000 - $84017 - $70011) * 16.67% = $58340